# Image Matching Filtering and Refinement by Planes and Beyond

### v3

@@ Abstract
This paper introduces a modular, non-deep learning method for filtering and refining sparse correspondences in image matching. Assuming that motion flow within the scene can be approximated by local homography transformations, matches are aggregated into overlapping clusters corresponding to virtual planes using an iterative RANSAC-based approach, with non-conforming correspondences discarded. Moreover, the underlying planar structural design provides an explicit map between local patches associated with the matches, enabling optional refinement of keypoint positions through cross-correlation template matching after patch reprojection. Finally, to enhance robustness and fault-tolerance against violations of the piece-wise planar approximation assumption, a further strategy is designed for minimizing relative patch distortion in the plane reprojection by introducing an intermediate homography that projects both patches into a common plane. The proposed method is extensively evaluated on standard datasets and image matching pipelines, and compared with state-of-the-art approaches. Unlike other current comparisons, the proposed benchmark also takes into account the more general, real, and practical cases where camera intrinsics are unavailable. Experimental results demonstrate that our proposed non-deep learning, geometry-based approach achieves performances that are either superior to or on par with recent state-of-the-art deep learning methods. Finally, this study suggests that there are still development potential in actual image matching solutions in the considered research direction, which could be in the future incorporated in novel deep image matching architectures.
@@
image matching, keypoint refinement, planar homography, normalized cross-correlation, SIFT, SuperGlue, LoFTR1 Introduction
@@ 1.1 Background
Image matching constitutes the backbone of most higher-level computer vision tasks and applications that require image registration to recover the 3D scene structure, including the camera poses. Image stitching [1, 2], Structure-from-Motion (SfM) [2, 3, 4, 5], Simultaneous Localization and Mapping (SLAM) [6, 2] and the more recently Neural Radiance Fields (NeRF) [7, 8] and Gaussian Splatting (GS) [9, 10] are probably nowadays the most common, widespread and major tasks critically relying on image matching.
Image matching traditional paradigm can be organized into a modular pipeline concerning keypoint extraction, feature description, and the proper correspondence matching [11]. This representation is being neglected in modern deep end-to-end image matching methods due to their intrinsic design which guarantees a better global optimization at the expense of the interpretability of the system. Notwithstanding that end-to-end deep networks represent for many aspects the State-Of-The-Art (SOTA) in image matching, such in terms of the robustness and accuracy of the estimated poses in complex scenes [12, 13, 14, 15, 16, 17, 18], handcrafted image matching pipelines such as the popular Scale Invariant Feature Transform (SIFT) [19] are employed still today in practical applications [4, 2] due to their scalability, adaptability, and understandability. Furthermore, handcrafted or deep modules remain present in post-process modern monolithic deep architectures, to further filter the final output matches as done notably by the RAndom SAmpling Consensus (RANSAC) [20] based on geometric constraints.
Image matching filters to prune correspondences are not limited to RANSAC, which in any of its form [21, 22, 23, 24] still remains mandatory nowadays as final step, and have been evolved from handcrafted methods [25, 26, 27, 28, 29, 30, 31, 32, 33] to deep ones [34, 35, 36, 37, 38, 39, 40, 41]. In addition, more recently methods inspired by coarse-to-fine paradigm, such as the Local Feature TRansformer (LoFTR) [14], have been developed to refine correspondences [42, 43, 44], which can be considered as the introduction of a feedback-system in the pipelined interpretation of the matching process.
@@ 1.2 Main contribution
As contribution, within the former background, this paper introduces a modular approach to filter and refine matches as post-process before RANSAC. The key idea is that the motion flow within the images, i.e. the correspondences, can be approximated by local overlapping homography transformations [45]. This is the core concept of traditional image matching pipelines, where local patches are approximated by less constrained transformations going from a fine to a coarse scale level. Specifically, besides the similarity and affine transformations, the planar homography is introduced at a further level. The whole approach proceeds as follows:
@@
The Multiple Overlapping Planes (MOP) module aggregates matches into soft clusters, each one related to a planar homography roughly representing the transformation of included matches. On one hand, matches unable to fit inside any cluster are discarded as outliers, on the other hand, the associated planar map can be used to normalize the patches of the related keypoint correspondences. As more matches are involved in the homography estimation, the patch normalization process becomes more robust than the canonical ones used for instance in SIFT. Furthermore, homographies better adapt the warp as they are more general than similarity or affine transformations. MOP is implemented by iterating RANSAC to greedily estimate the next best plane according to a multi-model fitting strategy. The main difference from previous similar approaches [46, 47], is in maintaining matches during the iterations in terms of strong and weak inliers according to two different reprojection errors to guide the cluster partitioning.
@@
The Middle Homography (MiHo) module additionally improves the patch normalization by further minimizing the relative patch distortion. More in detail, instead of reprojecting the patch of one image into the other image taken as reference according to the homography associated with the cluster, both the corresponding patches are reprojected into a virtual plane chosen to be in the middle of the base homography transformations so that globally the deformation is distributed into both the two patches, reducing the patch distortions with respect to the original images due to interpolation but also to the planar approximation.
@@
The Normalized Cross-Correlation (NCC) is employed following traditional template matching [48, 2] to improve the keypoint localization in the reference middle homography virtual plane using one patch as a template to being localized on the other patch. The relative shift adjustment found on the virtual plane is finally back-projected through into the original images. In order to improve the process, the template patch is perturbed by small rotations and anisotropic scaling changes for a better search in the solution space.
@@
MOP and MiHo are purely geometric-based, requiring only keypoint coordinates, while NCC relies on the image intensity in the local area of the patches. The early idea of MOP+MiHo+NCC can be found in [49]. Furthermore, MOP overlapping plane exploitation concept have been introduced in [50], while MiHo definition was draft in [51]. The proposed approach is handcrafted and hence more general as is does not require re-training to adapt to the particular kind of scene and its behavior is more interpretable as not hidden by deep architectures. As reported and discussed later in this paper, RANSAC takes noticeable benefits from MOP+MiHo as pre-filter before it in most configurations and scene kinds, and in any case does not deteriorate the matches. Concerning MOP+MiHo+NCC instead, keypoint localization accuracy seems to strongly depend on the kind of extracted keypoint. In particular, for corner-like keypoints, which are predominant in detectors such as the Keypoint Network (Key.Net) [52] or SuperPoint [53], NCC greatly improves the keypoint position, while for blob-like keypoints, predominant for instance in SIFT, NCC application can degrade the keypoint localization, due probably to the different nature of their surrounding areas. Nevertheless, the proposed approach indicates that there are still margins for improvements in current image matching solutions, even the deep-based ones, exploiting the analyzed methodology.
@@ 1.3 Additional contribution
As further contribution, an extensive comparative evaluation is provided against eleven less and more recent SOTA deep and handcrafted image matching filtering methods on both non-planar indoor and outdoor standard benchmark datasets [54, 55, 11], as well as planar scenes [56, 57]. Each filtering approach, including several combinations of the three proposed modules, has been tested to post-process matches obtained with seven different image matching pipelines and end-to-end networks, also considering distinct RANSAC configurations. Moreover, unlike the mainstream evaluation which has been taken as current standard [14], the proposed benchmark assumes that no camera intrinsics are available, reflecting a more general, realistic, and practical scenario. On this premise, the pose error accuracy is defined, analyzed, and employed with different protocols intended to highlight the specific complexity of the pose estimation working with fundamental and essential matrices [45], respectively. This analysis reveal the many-sided nature of the image matching problem and of its solutions.
@@ 1.5 Additional resources
Code and data are freely available online111https://github.com/fb82/MiHo. Data also include high-resolution versions of the images, plots, and tables presented in this paper222https://github.com/fb82/MiHo/tree/main/data/paper.
@@ 2.1 Disclamer and further references
The following discussion is far to be exhaustive due to more than three decades of active research and progress in the field, which is still ongoing. The reader can refer to [58, 59, 60, 61] for a more comprehensive analysis than that reported below, which is related and covers only the methods discussed in the rest of the paper.
@@ 2.2.1 Keypoint detectors
Keypoint detection is aimed at finding salient local image regions whose feature descriptor vectors are at the same time-invariant to image transformations, i.e. repeatable, and distinguishable from other non-corresponding descriptors, i.e. discriminant. Clearly, discriminability decreases as long as invariance increases and vice-versa. Keypoints are usually recognized as corner-like and blob-like ones, even if there is no clear separation within them outside ad-hoc ideal images. Corners are predominantly extracted by the Harris detector [62], while blobs by SIFT [19]. Several other handcrafted detectors exist besides the Harris and the SIFT one [63, 64, 65], but excluding deep image matching nowadays the former and their extensions are the most commonly adopted. Harris and SIFT keypoint extractors rely respectively on the covariance matrix of the first-order derivatives and the Hessian matrix of the second-order derivatives. Blobs and corners can be considered each other analogues for different derivative orders, as in the end, a keypoint in both cases correspond to a local maxima response to a filter which is a function of the determinant of the respective kind of matrix. Blob keypoint centers tend to be pushed away from edges, which characterize the regions, more than corners due to the usage of high-order derivatives. More recently, the joint extraction and refinement of robust keypoints has been investigated by the Gaussian Mixture Models for Interpretable Keypoint Refinement and Scoring (GMM-IKRS) [66]. The idea is to extract image keypoints after several homography warps and cluster them after back-projection into the input images so as to select the best clusters in terms of robustness and deviation, as another perspective of the Affine SIFT (ASIFT) [67] strategy.
@@ 2.2.2 Keypoint descriptors
Feature descriptors aim at embedding the keypoint characterization into numerical vectors. There has been a long research interest in this area, both in terms of handcrafted methods and non-deep machine learning ones, which contributed actively to the current development and the progress made by current SOTA deep image matching. Among non-deep keypoint descriptors, nowadays only the SIFT descriptor is practically still used and widespread due to its balance within computational efficiency, scaling, and matching robustness. SIFT descriptor is defined as the gradient orientation histogram of the local neighborhood of the keypoint, also referred to as patch. Many high-level computer vision applications like COLMAP [4] in the case of SfM are based on SIFT. Besides SIFT keypoint detector & descriptor, the Oriented FAST and Rotated BRIEF (ORB) [68] is a pipeline surviving still today in the deep era worth to be mentioned. ORB is the core of the popular ORB-SLAM [6], due to its computational efficiency, even if less robust than SIFT. Nevertheless, it seems to be going to be replaced by SIFT with the increased computational capability of current hardware [69].
@@ 2.2.3 Patch normalization
Actually, the patch normalization takes place in between the detection and description of the keypoints, and it is critical for effective computation of the descriptor to be associated with the keypoint. Patch normalization is implicitly hidden inside both keypoint extraction and description and its task roughly consists of aligning and registering patches before descriptor computation [19]. The key idea is that complex image deformations can be locally approximated by more simple transformations as long as the considered neighborhood is small. Under these assumptions, SIFT normalizes the patch by a similarity transformation, where the scale is inferred by the detector and the dominant rotation is computed from the gradient orientation histogram. Affine covariant patch normalization [70, 67, 71] was indeed the natural evolution, which can improve the normalization in case of high degrees of perspective distortions. Notice that in the case of only small scene rotation, up-right patches computed neglecting orientation alignment can lead to better matches since a strong geometric constraint limiting the solution space is imposed [72]. Concerning instead patch photometric consistency, affine normalization, i.e. by mean and standard deviation, is usually the preferred choice being sufficiently robust in the general case [19].
@@ 2.2.4 Descriptor similarity and global geometric constraints
The last step is the proper matching, which mainly pairs keypoints by considering the distance similarity on the associated descriptors. Nearest Neighbor (NN) and Nearest Neighbor Ratio (NNR) [19] are the common and simple approaches even if additional extensions have been developed [73, 32]. For many tasks descriptor similarity alone does not provide satisfactory results, which can be achieved instead by including a further filtering post-process by constraining matches geometrically. RANSAC provides an effective method to discard outliers under the more general epipolar geometry constraints of the scene or in the case of planar homographies. The standard RANSAC continues to be extended in several ways [21, 22]. Worth mentioning are the Degenerate SAmpling Consensus (DegenSAC) [24] to avoid degenerate configurations, and the MArGinalized SAmpling Consensus (MAGSAC) [23] for defining a robust inlier threshold which remains the most critical parameter.
@@ 2.2.5 Local geometric constraints
RANSAC exploits global geometric constraints. A more general and relaxed geometric assumption than RANSAC is to consider only local neighborhood consistency across the images as done in Grid-based Motion Statistics (GMS) [26] using square neighborhood, but also circular [25] and affine-based [27, 28] neighborhoods or Delaunay triangulation [32, 28] have been employed. In most cases, these approaches require an initial set of robust seed matches for initialization whose selection may become critical [74], while other approaches explicitly exploit descriptor similarities to rank matches [32]. In this respect, the vanilla RANSAC and GMS are pure geometric image filters. Notice that descriptor similarities have been also exploited successfully within RANSAC [75]. Closely connected to local filters, another class of methods is designed instead to estimate and interpolate the motion field upon which to check match consistency [29, 31, 30]. Local neighborhood consistency can be framed also into RANSAC as done by Adaptive Locally-Affine Matching (AdaLAM) [33] which runs multiple local affine RANSACs inside the circular neighborhoods of seed matches. AdaLAM design can be related with GroupSAC [76], which draws hypothesis model sample correspondences from different clusters, and with the more general multi-model fitting problem for which several solutions have been investigated [77, 78]. Among these approaches the recent Consensus Clustering (CC) [79] has been developed for the specific case of multiple planar homographies.

### v4

@@ Abstract
This paper introduces a modular, non-deep learning method for filtering and refining sparse correspondences in image matching. Assuming that motion flow within the scene can be approximated by local homography transformations, matches are aggregated into overlapping clusters corresponding to virtual planes using an iterative RANSAC-based approach discarding incompatible correspondences. Moreover, the underlying planar structural design provides an explicit map between local patches associated with the matches, by which optionally refine the keypoint positions through cross-correlation template matching after the patch reprojection. Finally, to enhance robustness and fault-tolerance against violations of the piece-wise planar approximation assumption, a further strategy is designed in order to minimize the relative patch distortion in the plane reprojection by introducing an intermediate homography that projects both patches into a common plane. The proposed method is extensively evaluated on standard datasets and image matching pipelines, and compared with state-of-the-art approaches. Unlike other current comparisons, the proposed benchmark also takes into account the more general, real, and practical cases where camera intrinsics are unavailable. Experimental results demonstrate that our proposed non-deep learning, geometry-based filter is effective in presence of outliers and the optional cross-correlation refinement step is valid in the case of corner-like keypoints. Finally, this study suggests that there is still significant development potential in practical image matching solutions in the considered research direction, which could be in the future incorporated in novel deep image matching architectures.
@@
image matching, keypoint refinement, planar homography, normalized cross-correlation, SIFT, SuperGlue, LoFTR, MASt3R, RoMa1 Introduction
@@ 1.1 Background
Image matching constitutes the backbone of most higher-level computer vision tasks and applications that require image registration to recover the 3D scene structure, including the camera poses. Image stitching [stitching, zeliski_book], Structure-from-Motion (SfM) [zeliski_book, phototourism, colmap, glomap], Simultaneous Localization and Mapping (SLAM) [orbslam, zeliski_book] and the more recently Neural Radiance Fields (NeRF) [nerf, nerf_survey] and Gaussian Splatting (GS) [gaussian_splatting, gaussian_splatting_survey] are probably nowadays the most common, widespread and major tasks critically relying on image matching.
Image matching traditional paradigm can be organized into a modular pipeline concerning keypoint extraction, feature description, and the proper correspondence matching [imw2020]. This representation is being neglected in modern deep end-to-end image matching methods due to their intrinsic design which guarantees a better global optimization at the expense of the interpretability of the system. Notwithstanding that end-to-end deep networks represent for many aspects the State-Of-The-Art (SOTA) in image matching, such in terms of the robustness and accuracy of the estimated poses in complex scenes [lightglue, aliked, loftr, quadtree, matchformer, dust3r, roma], handcrafted image matching pipelines such as the popular Scale Invariant Feature Transform (SIFT) [sift] are employed still today in practical applications [colmap, zeliski_book] due to their scalability, adaptability, and understandability. Furthermore, handcrafted or deep modules remain present as match post-processing, in order to further filter the final output matches as done notably by the RAndom SAmpling Consensus (RANSAC) [ransac], based on geometric constraints.
Image matching filters to prune correspondences are not limited to RANSAC, which in any of its form [usac, gcransac, magsac, degensac] still remains mandatory nowadays as final step, and have been evolved from handcrafted methods [lpm, gms, pgm, pfm, vfc, code, bm, dtm, adalam] to deep ones [learning_correspondences, acne, oanet, clnet, ncmnet, ms2dgnet, convmatch, dematch]. In addition, more recently methods inspired by coarse-to-fine paradigm, such as the Local Feature TRansformer (LoFTR) [loftr], have been developed to refine correspondences [patch2pix, keypt2subpx, fcgnn], which can be considered as the introduction of a feedback-system in the pipelined interpretation of the matching process.
@@ 1.2 Main contribution
As contribution, within the former background, this paper introduces a modular approach to filter and refine matches as post-process before RANSAC. The key idea is that the motion flow within the images, i.e. the correspondences, can be approximated by local overlapping homography transformations [multiview]. This is the core concept of traditional image matching pipelines, where local patches are approximated by less constrained transformations going from a fine to a coarse scale level. Specifically, besides the similarity and affine transformations, the planar homography is introduced at a further level. The whole approach proceeds as follows:
@@
The Multiple Overlapping Planes (MOP) module aggregates matches into soft clusters, each one related to a planar homography roughly representing the transformation of the included matches. On one hand, matches unable to fit inside any cluster are discarded as outliers; on the other hand, the associated planar map can be used to normalize the patches of the related keypoint correspondences. As more matches are involved in the homography estimation, the patch normalization process becomes more robust than the canonical ones used for instance in SIFT. Furthermore, homographies better adapt the warp as they are more general than similarity or affine transformations. MOP is implemented by iterating RANSAC to greedily estimate the next best plane according to a multi-model fitting strategy. The main difference from previous similar approaches [seqransac1, seqransac2], is in maintaining matches during the iterations in terms of strong and weak inliers according to two different reprojection errors to guide the cluster partitioning.
@@
The Middle Homography (MiHo) module additionally improves the patch normalization by further minimizing the relative patch distortion. More in detail, instead of reprojecting the patch of one image into the other image taken as reference according to the homography associated with the cluster, both the corresponding patches are reprojected into a virtual plane chosen to be in the middle of the base homography transformations. The globally deformation gets distributed into both the two patches, reducing the patch distortions with respect to the original images due to interpolation but also to the planar approximation. With respect to the half homography described in [VSAC], MiHo is more relaxed in adhering to the planar constraints and introduces less distortions in the regions of the images which do not overlap.
@@
The Normalized Cross-Correlation (NCC) is employed following traditional template matching [gonzales_wood, zeliski_book] to improve the keypoint localization in the reference middle homography virtual plane using one patch as a template to being localized on the other patch. The relative shift adjustment found on the virtual plane is finally back-projected through into the original images. In order to improve the process, the template patch is perturbed by small rotations and anisotropic scaling changes for a better search in the solution space. NCC search space can be further limited according to the patch gradient autocorrelation, improving the refinement.
@@
MOP and MiHo are purely geometric, requiring only keypoint coordinates, while NCC relies on the image intensity in the local area of the patches. The early idea of MOP+MiHo+NCC can be found in [miho_base]. Furthermore, MOP overlapping plane exploitation concept have been introduced in [slime], while MiHo definition was draft in [average_homography]. The proposed approach is handcrafted and hence more general as is does not require re-training to adapt to the particular kind of scene and its behavior is more interpretable as not hidden by deep architectures. As reported and discussed later in this paper, RANSAC takes noticeable benefits from MOP+MiHo as pre-filter before it in most configurations and scenes when the outlier contamination of the base pipeline is relevant, and in any case does not deteriorate the matches.
Concerning MOP+MiHo+NCC, keypoint localization accuracy seems to strongly depend on the kind of extracted keypoint. In particular, for corner-like keypoints, which are predominant in detectors such as the Keypoint Network (Key.Net) [keynet] or SuperPoint [superpoint], NCC greatly improves the keypoint position. For blob-like keypoints, predominant for instance in SIFT, or flat patches, NCC can degrade the keypoint localization, due probably to the different nature of their surrounding areas. Nevertheless, the proposed approach indicates that there are still margins for improvements in current image matching solutions, even the deep-based ones, exploiting the analyzed methodology.
@@ 1.3 Additional contribution
As further contribution, an extensive comparative evaluation is provided against thirteen less and more recent SOTA deep and handcrafted image matching filtering and refinement methods. The evaluation includes both non-planar indoor and outdoor standard benchmark datasets [megadepth, scannet, imw2020], as well as planar scenes [hpatches, mods]. Each filtering approach, including several combinations of the three proposed modules, has been tested to post-process matches obtained with nine different image matching pipelines and end-to-end networks, also considering distinct RANSAC configurations. Moreover, unlike the mainstream evaluation which has been taken as current standard [loftr], the proposed benchmark assumes that no camera intrinsics are available, reflecting a more general, realistic, and practical scenario. On this premise, the pose error accuracy is defined, analyzed, and employed with different protocols intended to highlight the specific complexity of the pose estimation working with fundamental and essential matrices [multiview]. This analysis reveals the many-sided nature of the image matching problem and of its solutions.
@@ 1.5 Additional resources
Code and data are freely available online111https://github.com/fb82/MiHo. Data also include further experimental evaluations, high-resolution versions of the images, plots, and tables presented in this paper222https://github.com/fb82/MiHo/tree/main/data/paper.
@@ 2.1 Disclamer and further references
The following discussion is far to be exhaustive due to more than three decades of active research and progress in the field, which is still ongoing. The reader can refer to [im_survey, im_survey_other, im_survey_another, im_survey_yet_another] for a more comprehensive analysis than that reported below, which is related and covers only the methods discussed in the rest of the paper.
@@ 2.2.1 Keypoint detectors
Keypoint detection is aimed at finding salient local image regions whose feature descriptor vectors are at the same time-invariant to image transformations, i.e. repeatable, and distinguishable from other non-corresponding descriptors, i.e. discriminant. Clearly, discriminability decreases as long as invariance increases and vice-versa. Keypoints are usually recognized as corner-like and blob-like ones, even if there is no clear separation within them outside ad-hoc ideal images. Corners are predominantly extracted by the Harris detector [harris], while blobs by SIFT [sift]. Several other handcrafted detectors exist besides the Harris and the SIFT ones [akaze, surf, fast], but excluding deep image matching nowadays the former and their extensions are the most commonly adopted. Harris and SIFT keypoint extractors rely respectively on the covariance matrix of the first-order derivatives and the Hessian matrix of the second-order derivatives. Blobs and corners can be considered each other analogues for different derivative orders, as in the end, a keypoint in both cases correspond to a local maxima response to a filter which is a function of the determinant of the respective kind of matrix. Blob keypoint centers tend to be pushed away from edges, which characterize the regions, more than corners due to the usage of high-order derivatives. More recently, the joint extraction and refinement of robust keypoints has been investigated by the Gaussian Mixture Models for Interpretable Keypoint Refinement and Scoring (GMM-IKRS) [gmmikrs]. The idea is to extract image keypoints after several homography warps and cluster them after back-projection into the input images so as to select the best clusters in terms of robustness and deviation, as another perspective of the Affine SIFT (ASIFT) [asift] strategy.
@@ 2.2.2 Keypoint descriptors
Feature descriptors aim at embedding the keypoint characterization into numerical vectors. There has been a long research interest in this area, both in terms of handcrafted methods and non-deep machine learning ones, which contributed actively to the current development and the progress made by current SOTA deep image matching. Among non-deep keypoint descriptors, nowadays only the SIFT descriptor is practically still used and widespread due to its balance within computational efficiency, scaling, and matching robustness. SIFT descriptor is defined as the gradient orientation histogram of the local neighborhood of the keypoint, also referred to as patch. Many high-level computer vision applications, like COLMAP [colmap] in the case of SfM, are based on SIFT. Besides SIFT keypoint detector & descriptor, the Oriented FAST and Rotated BRIEF (ORB) [orb] is a pipeline surviving still today in the deep era worth to be mentioned. ORB is the core of the popular ORB-SLAM [orbslam], due to its computational efficiency, even if less robust than SIFT. Nevertheless, it seems to be going to be replaced by SIFT with the increased computational capability of current hardware [cudasift].
@@ 2.2.3 Patch normalization
Actually, the patch normalization takes place in between the detection and description of the keypoints, and it is critical for effective computation of the descriptor to be associated with the keypoint. Patch normalization is implicitly hidden inside both keypoint extraction and description, and its task roughly consists of aligning and registering patches before the descriptor computation [sift]. The key idea is that complex image deformations can be locally approximated by more simple transformations as long as the considered neighborhood is small. Under these assumptions, SIFT normalizes the patch by a similarity transformation, where the scale is inferred by the detector and the dominant rotation is computed from the gradient orientation histogram. Affine covariant patch normalization [hess_lapl_affine, asift, sift_affine] was indeed the natural evolution, which can improve the normalization in case of high degrees of perspective distortions. Notice that in the case of only small scene rotation, up-right patches computed neglecting orientation alignment can lead to better matches since a strong geometric constraint limiting the solution space is imposed [sgloh2]. Concerning instead patch photometric consistency, affine normalization, i.e. by mean and standard deviation, is usually the preferred choice being sufficiently robust in the general case [sift].
@@ 2.2.4 Descriptor similarity and global geometric constraints
The last step is the proper matching, which mainly pairs keypoints by considering the distance similarity on the associated descriptors. Nearest Neighbor (NN) and Nearest Neighbor Ratio (NNR) [sift] are the common and simple approaches even if additional extensions have been developed [fginn, dtm]. For many tasks descriptor similarity alone does not provide satisfactory results, which can be achieved instead by including a further filtering post-process by constraining matches geometrically. RANSAC provides an effective method to discard outliers under the more general epipolar geometry constraints of the scene or in the case of planar homographies. The standard RANSAC continues to be extended in several ways [usac, gcransac, superransac]. Worth mentioning are the Degenerate SAmpling Consensus (DegenSAC) [degensac] to avoid degenerate configurations, the MArGinalized SAmpling Consensus (MAGSAC) [magsac] for defining a robust inlier threshold, which remains the most critical parameter, and the Verified SAmpling Consensus (VSAC) [VSAC] which checks and penalizes outliers that can appear as inliers merely by chance.
@@ 2.2.5 Local geometric constraints
RANSAC exploits global geometric constraints. A more general and relaxed geometric assumption than RANSAC is to consider only local neighborhood consistency across the images, as done in Grid-based Motion Statistics (GMS) [gms] using square neighborhood. Circular [lpm] and affine-based [pgm, pfm] neighborhoods, or Delaunay triangulation [dtm, pfm], have been employed too. In most cases, these approaches require an initial set of robust seed matches for initialization, whose selection may become critical [glpm], while other approaches explicitly exploit descriptor similarities to rank matches [dtm]. In this respect, the vanilla RANSAC and GMS are pure geometric image filters. Notice that descriptor similarities have been also exploited successfully within RANSAC [prosac]. Closely connected to local filters, another class of methods is designed instead to estimate and interpolate the motion field upon which to check match consistency [vfc, bm, code]. Local neighborhood consistency can be framed also into RANSAC as done by Adaptive Locally-Affine Matching (AdaLAM) [adalam], which runs multiple local affine RANSACs inside the circular neighborhoods of seed matches. AdaLAM design can be related with GroupSAC [groupsac], which draws hypothesis model sample correspondences from different clusters, and with the more general multi-model fitting problem for which several solutions have been investigated [jlinkage]. Among these approaches Consensus Clustering (CC) [cc] and Progressive-X [progx] have been developed for the specific case of multiple planar homographies. Both CC and Progressive-X are designed for extracting strict planes and rely on tight reprojection error thresholds, in contrast to the proposed MOP-based approaches that relies on a more relaxed design, and hence is able to work with approximated virtual and real planes of the scene.